# Approximation in Lattice Field Theories

### v1

@@ Abstract
We investigate an approximation to the Schwinger-Dyson (SD) equations of the collective Coulomb field of the large $N$ homogeneous electron fluid. The large $N$ approximation transforms the infinite SD hierarchy is into a set of closed, equations for 1 and 2-pt correlators. In this paper, the dynamics of a toy model - a small square Euclidean lattice with periodic boundary conditions - are considered. The Markov Chain Monte Carlo numerical method evaluated the 1 and 2-pt correlation functions on a $2\times 2$ and $3\times 3$ lattice. The derived equations are checked with the correlator values, and an agreement at $N\sim 10^{3}$ to order $10^{-3}$ was found. The agreement can be further strengthened by increasing runs in the Markov Chain Monte Carlo method.
@@ 1 Introduction
In [1], one of us presented a method for re-summing perturbative and vector $1/N$ expansions by using Schwinger-Dyson (SD) equations. The basic problem with SD equations is that they form an infinite hierarchy. The observation of [1] was that, when written in terms of one particle irreducible (1PI) correlation functions, in a model with a small coupling, this hierarchy could be systematically truncated by noting that the $k$ point 1PI correlator vanishes like $g^{k}$. Each level of truncation gives rise to a finite closed system of non-linear integral equations, which are amenable to solution on modern computers. The same methods apply to the SD equations for the collective fields of vector $1/N$ expansions, where they are a generalization and improvement of a technique proposed in [2].
@@
The paper is structured as follows. In Section 2, we derive the truncation of the infinite hierarchical equations up to order $1/N$. Then, we describe the numerical method needed to evaluate the correlators in Section 3. We also thoroughly explain the specifics of the numerical technique when applied to our problem. Finally, the results of the code, along with the code, are presented in Section 4. We find agreement at the $10^{-3}$ accuracy level for $N\sim 10^{3}$ on the $2\times 2$ lattice and on the $3\times 3$ lattice. While this accuracy is already strong, we also explain how to increase it.
@@ 2 Derivation of Closed Equations
Throughout this manuscript, $A_{ij}\equiv K_{ij}+\Phi_{0}(i)\mathbb{1}_{ij}$, $K_{ij}$ is the $M\times M$ 2D discrete Laplacian matrix with periodic boundary conditions, and $(A^{-1})_{ij}$ is the propagator between lattice sites $i$ and $j$. We can then expand the action perturbatively, since we are interested in the terms that dominate in the action for large $N$. Ignoring the constant and linear terms, the action takes the form:
@@
where $S^{(2)}(i,j)=\delta_{ij}-(A^{-1})_{ij}(A^{-1})_{ji}$ and $S^{(k)}(i_{1},\cdots,i_{k})=(-1)^{k}(k-1)!(A^{-1})_{i_{1}i_{2}}(A^{-1})_{i_{2}i_{3}}\cdots(A^{-1})_{i_{k}i_{1}}$. The perturbative expansion provides a straightforward method to determine the stationary point, which we will cover in detail in Section 4. Figure 2 contains the Feynman diagrams for $S^{(2)},S^{(3)},$ and $S^{(4)}$. Diagrams for higher order terms simply follow this pattern.
@@
$\begin{split}j(i)=&\sum_{k}^{M}S^{(2)}(i,k)W_{k}+\sum_{k,l}^{M}\frac{1}{2\sqrt{N}}S^{(3)}(i,k,l)(W_{k}W_{l}-W_{kl})\\ &-\sum_{k,l,m}^{M}\frac{1}{6N}S^{(4)}(i,k,l,m)\bigg{[}W_{kl}W_{m}+W_{km}W_{l}+W_{lm}W_{k}-W_{klm}-W_{k}W_{l}W_{m}\bigg{]}\end{split}$ (4)
@@
$\begin{split}0=&\sum_{k}^{M}S^{(2)}(i,k)W_{kt}+\frac{1}{2\sqrt{N}}S^{(3)}(i,k,l)\bigg{[}W_{kt}W_{l}+W_{k}W_{lt}-W_{klt}\bigg{]}-\delta_{it}\\ &-\frac{1}{6N}S^{(4)}(i,k,l,m)\bigg{[}(W_{klt}W_{m}+W_{kl}W_{mt}+\text{symmetric terms})-W_{klmt}-(W_{kt}W_{l}W_{m}+\text{symmetric terms})\bigg{]}.\end{split}$ (5)
Only $M(M+1)/2$ out of the $M^{2}$ equations are unique, since derivatives commute. Eqns. (4) and (5), which consist of $M+M(M+1)/2$ equations in total, serve as the metrics for how successful this approximation is. Note that the sources $j(i)$ are set to 0 in eqn. (4) by definition.
@@ 2.1 Approximating Higher Order Correlation Functions
By definition, we have that $W_{i}$ is the 1-pt correlation function $\langle x(i)\rangle$. There are only $M$ such terms, so these terms are manageable. Similarly, calculating $\langle x(i)x(j)\rangle$ gives us $W_{ij}$. However, working with 3 and 4-pt correlation functions is a lot more computationally intensive, especially given the large number of degrees of freedom.
[…]
@@
$\begin{split}W_{3}(k,l,m)\equiv W_{klm}&=\frac{\partial^{3}W}{\partial j(k)\partial j(l)\partial j(m)}=-\frac{\partial}{\partial j(m)}(\Gamma_{kl}^{-1})=-\sum_{a}^{M}\frac{\partial\phi(a)}{\partial j(m)}\frac{\partial}{\partial\phi(a)}(\Gamma_{kl}^{-1})=\sum_{a,b,c}^{M}W_{am}W_{bk}\Gamma_{abc}W_{cl}\\ &\approx\sum_{a,b,c}^{M}W_{am}W_{bk}S^{(3)}(a,b,c)W_{cl}.\end{split}$ (7)
Above, $\Gamma^{-1}_{kl}=(\partial^{2}\Gamma/\partial s_{c}(k)\partial s_{c}(l))^{-1}$ - a fact that directly follows from eqn. (6). Similarly, $W_{klmn}$ can be expressed as:
$\begin{split}W_{4}(k,l,m,n)\equiv W_{klmn}=&\sum_{a,b,c}^{M}\bigg{[}W_{amn}W_{bk}S^{(3)}(a,b,c)W_{cl}+W_{bkn}W_{am}S^{(3)}(a,b,c)W_{cl}+W_{cnl}W_{bk}S^{(3)}(a,b,c)W_{am}\bigg{]}\\ +&\sum_{a,b,c,d}^{M}W_{dn}W_{am}W_{bk}S^{(4)}(a,b,c,d)W_{cl}.\end{split}$ (8)
A tree diagram representation of eqns. (7) and (8) can be found in Figures 3 and 4. In the derivations up until now, one might have argued that one could ignore the $1/N$ term, since we are considering the large $N$ limit. However, we included $W_{4}$ to account for its potentially large contribution given its complex structure.
@@
Figure 4: A Tree diagram illustrating the relation between $W_{4}$ and $\Gamma_{4}$. Note that the $W_{3}$ term that appears in eqn. (8) has already been solved for in terms of $\Gamma_{3}$. For large enough NN, $\Gamma_{4}=S^{(4)}$.
With these substitutions, eqns. (4) and (5) can be expressed entirely in terms of $W_{i}$ and $W_{ij}$ , drastically reducing the number of unknowns in our equations to $M+M(M+1)/2$. Since this is also the exact number of equations, either the system has one solution or none. We attempted to utilize Solve and NSolve on Mathematica to explicitly find a solution (or lack thereof). However, such a code was unable to finish running in a reasonable time, even when considering $M=4$ and ignoring the $1/N$ terms. Thus, we turned to a numerical investigation of the 1 and 2-pt connected correlation functions. If the numerically evaluated values satisfy the closed equations, then the approximation is successful.
@@ 3 Specifics of the Numerical Analysis
The connected correlation functions cannot be explicitly calculated, since $Z_{0}$ and its respective numerator diverge. Even if it were possible, such a calculation would be unreasonable for higher dimensional integrals (i.e. $M\gg 1$). This is especially troublesome for us, given that $M$ scales as a quadratic as the Euclidean lattice is taken to be a square.
@@ 3.1 Introduction to Markov Chain Monte Carlo
Markov Chain Mote Carlo (MCMC) is a numerical method that explicitly combats both of these issues [3, 4]. The fundamental strategy of this technique is to construct a Markov Chain, the mean of which is the desired value, for each observable. A simple uniform probability distribution will also yield accurate results, but the variance of the elements would fall too slowly, especially when working with many observables [4]. The MCMC method reduces variance by introducing two different ideas: (i) rather than uniformly choosing integration variable values, it chooses them based on a probability distribution - this is a technique called importance sampling - and (ii) element $N$ of the Markov Chain is chosen based on only element N1N-1 and through importance sampling. Incorporating both of these properties reduces how fast the variance shrinks. The mathematics that validate this numerical technique, along with error analysis, can be found in [4].
@@
Above, we used the fact that $\int[dx]e^{-S}/Z_{0}=1$, which makes $e^{-S}/Z_{0}$ a valid probability distribution $\pi(x)$ [^1] We choose to use the same variable names as [4] throughout this manuscript to avoid confusion for readers who would like to expand their understanding of the numerical technique. for our purposes. Note that we can avoid evaluating $Z_{0}$ altogether to show this. Similarly, $W_{ij}=\langle x(i)x(j)\rangle$.
A concrete methodology to effectively utilize MCMC is given by the Metropolis-Hastings algorithm. While this is certainly not the only algorithm, it allows us to construct an irreducible, aperiodic Markov Chain with positive recurrent states - conditions necessary for the success of MCMC [4]. This algorithm revolves around updating the values of the action’s variable(s) (of a site) with the given probability distribution. If the action is multi variable - such as ours - then it is acceptable to either randomly choose a site to update or sequentially update each site. As stated prior, the updating process only takes in the current value $x(i)$ of a site to determine if the site can harbor the new value $\tilde{x}(i)$. That is, $P(\tilde{x}\leftarrow x)$ is only a function of $x,\tilde{x}$, and the probability distribution.
The general outline of the algorithm is as follows [3, 4]:
@@ 3.2 Applying Metropolis-Hasting
Before we can work with the algorithm, we must numerically compute two quantities in our action: the discrete Laplacian and the stationary point. The perturbative expansion of the action with respect to $\sqrt{N}$ tells us that $\Phi_{0}-(A^{-1})_{ii}=0$ for fixed $i$. This generally needs to be solved numerically, as $A_{ij}$ is a function of $\Phi_{0}$. The 2D discrete Laplacian with periodic boundary conditions can be explicitly calculated using a five-point stencil on a grid like in Figure 1, but this may be tedious for larger grids. A much easier alternative is using the Kronecker sum of two 1 dimensional discrete Laplacians $L_{1}$ with periodic boundary conditions. Thus, the discrete Laplacian $K_{ij}$ can be written as:
@@
Above, $\otimes$ represents a Kronecker product. All matrices on the right hand side above are $\sqrt{M}\times\sqrt{M}$ dimensional. This forces the 2D discrete Laplacian to be an $M\times M$ matrix. When constructed like this, $K_{ij}$ may not look translation invariant. That is, $K_{ij}\neq K(i-j)$. However, a simple rearranging of rows and columns can achieve this. Doing so does not change its spectrum, so the dynamics governed by the system can be studied using $K_{ij}$ from eqn. (11). There is also ambiguity in the sign of this matrix, based on the stencil chosen for $L_{1}$, but changing the sign simply adds a negligible constant in the action. The stationary point also changes sign, which does not affect any calculations.
An issue with using this calculated $K_{ij}$ is the inclusion of periodic boundary conditions. This forces each row to sum to 0, making the matrix singular. When solving $\Phi_{0}-(A^{-1})_{ii}=0$, we do not find valid solutions. Some of them are complex, while the rest are not local minimums of the action (when fluctuations are turned off) [^2] This can be checked by plotting the action (with fluctuations turned off) as a function of $\Phi_{0}$.. The lack of a stationary/equilibrium phase suggests that there is no stationary point here. To combat this, we redefined $K_{ij}$ by adding an identity matrix to it. Any matrix of the form $c\mathbb{1}$ would suffice, so we chose $c=1$ for simplicity. This rid the system of the singularity, allowing for a valid stationary point to exist. This extra term that we included can be thought of as a potential term.
@@
The action that we will use in the program is the closed-form variation so that we can avoid the infinite sum. However, this introduces a subtle source of error that has big ramifications. The two versions of the action are analytically the same except for two terms: the constant term and the linear term (with respect to $x(i)$). It is evident from eqn. (12) that the constant terms do not matter. One may be tempted to say the same thing about the linear term, but this is not true numerically. Choosing a stationary point that makes the linear term vanish is numerically equivalent to having an extremely small (and non-zero) term. Since the linear term is magnified by the corresponding factor of $\sqrt{N}$ attached to it, the desired integrals actually grow as $\sqrt{N}$. This can be avoided if we explicitly subtract out the linear term from the action. To potentially avoid unforeseen errors, we also subtracted out the constant terms.
While the action determines the likelihood of accepting a new value at a site, it does not play a role in proposing a new value at a site. There are multiple ways to do so, but in this version of the Metropolis-Hastings algorithm, we choose a point near the current value. That is, $\tilde{x}(j)=x(j)+\delta$, where $\delta$ is a uniformly and randomly chosen from an interval $\Delta,\Delta]$. There is no quantitative formula for $\Delta$, but the consensus is to choose $\Delta$ such that the program accepts new values at sites $50\%-60\%$ of the time. This choice is to maximize efficiency and precision [4].
@@
Figure 5: Plot of $\langle x_{3}\rangle$ for $M=4$ and $N=10^{2}$. All $x(i)$ of each run started with the same values, which were 0 (blue/solid), 2-2 (red/dashed), and 44 (black/dot-dashed) respectively. Note that after around $\sim 300$ updates, thermalization has been achieved.
The next phase is straightforward: we continuously conduct sweeps. Here, there are two simple, computational differences: (i) all previous sets of values must be stored and (ii) after each update, the observables calculated from the site values must be stored. However, only a subset of the observables’ arrays will be elements of the Markov Chain. Before we discuss how to determine this subset, we must understand why all the measurements cannot be considered. By definition of a stochastic process, all measurements are dependent on the directly prior set of measurements. Since the updates are typically small, the statistical dependence that one set of measurements shares with the next few sets of measurements will bar the MCMC error from decreasing [3]. For example, consider the following integral:
@@
The autocorrelation of the measurements can be utilized to discard these unnecessary measurements. If we were to apply Metropolis-Hasting to integrate $\int dx~f(x)p(x)$ and obtain measurements (not Markov Chain) $Y=\{f(y_{k})\}_{k=1}^{n}$, then the autocorrelation between $y_{i}$ and $y_{i+h}$ is defined as:
@@
When $\rho(h)$ is small enough, we can simply skip $h$ values in between measurements to construct our Markov Chain. There is not a consensus on this threshold, but we will use $\rho(h)<0.1$ [3, 4]. When the statistical dependence of the elements is this small, the error can simply be taken to be the Var($Y_{\text{MC}}$)/$n$, where $n$ is the length of the Markov Chain $Y_{\text{MC}}$. Thus, as $n\rightarrow\infty$, the exact value of the integral is reached.
Figure 6: Plot of $\rho(h)$ with respect to $h$ for the measurements of $\langle x^{2}(4)\rangle$ for $M=4$ and $N=10^{6}$.
To obtain these values in code, we update each site 10000 to obtain a large-enough array of measurements. While 10000 is an arbitrary number, a large-enough set of measurements is needed to ensure normal behavior from the variance. Then, we just keep calculating $\rho(h)$ until the first $h_{*}$ such that $\rho(h)<0.1$. An example of what these typically behave like can be found in Figure 6. Each site $x(i)$ may have a slightly different $h_{*}$ [^3] As far as we can tell, the length of the Markov Chain does not affect $h_{*}$ drastically.. We then continue the measurement phase and construct our Markov Chains with every $h_{*}$ measurements for each site. Figure 7 illustrates $\langle x^{2}(1)\rangle$, from different numbers of measurements, along with their respective errors as an example.

### v2

# Numerical Verification of Perturbative Schwinger-Dyson Resummation on Lattice Models

@@ Abstract
We investigate an approximation to the Schwinger-Dyson (SD) equations of the collective Coulomb field of the large $N$ Homogeneous Electron Fluid. The large $N$ approximation transforms the infinite SD hierarchy into a set of closed, equations for 1 and 2-pt correlators. In this paper, the dynamics of a toy model — a small, square Euclidean lattice with periodic boundary conditions — are considered. The Markov Chain Monte Carlo numerical method evaluated the 1 and 2-pt correlation functions on a $2\times 2$ and $3\times 3$ lattice. The derived equations are checked with the correlator values, and an agreement at $N\sim 100$ to order $10^{-3}$ was found. The agreement can be further strengthened by increasing runs in the Markov Chain Monte Carlo method.
Keywords:
Boundary Quantum Field Theory, Correlation Functions, Algorithms and Theoretical Development
arxiv: 2509.08501
@@ 1 Introduction
In tb1, one of us presented a method for re-summing perturbative and vector $1/N$ expansions by using Schwinger-Dyson (SD) equations. The basic problem with SD equations is that they form an infinite hierarchy. The observation of tb1 was that, when written in terms of one particle irreducible (1PI) correlation functions, in a model with a small coupling, this hierarchy could be systematically truncated by noting that the $k$ point 1PI correlator vanishes like $g^{k}$. Each level of truncation gives rise to a finite closed system of non-linear integral equations, which are amenable to solution on modern computers. The same methods apply to the SD equations for the collective fields of vector $1/N$ expansions, where they are a generalization and improvement of a technique proposed in banks2020systematic.
@@
The paper is structured as follows. In section 2, we derive the truncation of the infinite hierarchical equations up to order $1/N$. Then, we describe the numerical method needed to evaluate the correlators in section 3. We also thoroughly explain the specifics of the numerical technique when applied to our problem. Finally, the results of the code, along with the code, are presented in section 4. We find agreement at the $10^{-3}$ accuracy level for $N\sim 100$ on the $2\times 2$ lattice and on the $3\times 3$ lattice. While this accuracy is already strong, we also explain how to increase it.
@@ 2 Derivation of Closed Equations
Throughout this manuscript, $A_{ij}\equiv K_{ij}+\Phi_{0}\mathbb{1}_{ij}$, $K_{ij}$ is the $M\times M$ 2D discrete Laplacian matrix with periodic boundary conditions, and $(A^{-1})_{ij}$ is the propagator between lattice sites $i$ and $j$. With appropriate choices for the matrix $K_{ij}$ this model could approximate a large number of continuum bosonic field theories, with either relativistic or non-relativistic kinematics for the bosons and repulsive point-like interactions. The determinant comes from integrating out $N$ identical bosonic fields. If we changed the sign of the logarithm it would instead describe fermionic models like the large $N$ Hubbard model or the lattice version of the Gross-Neveu model. We could also replace the ultra-local quadratic term of the $x(i)$ variables with a lattice Laplacian or Helmholtz operator, to describe particles interacting via Coulomb or Yukawa forces. In this paper, since we are just testing the numerical accuracy of our proposed approximation method, we will restrict attention to bosons and choose $K_{ij}$ to be a lattice approximation to a two dimensional Laplacian. Thus, we will be studying the large $N$ relativistic $(\vec{\phi}^{2})^{2}$ model on a very small two dimensional lattice.
We can then expand the action perturbatively, since we are interested in the terms that dominate in the action for large $N$. Ignoring the constant and linear terms, the action takes the form:
@@
where $S^{(2)}(i,j)=\delta_{ij}-(A^{-1})_{ij}(A^{-1})_{ji}$ and $S^{(k)}(i_{1},\cdots,i_{k})=(-1)^{k-1}(k-1)!(A^{-1})_{i_{1}i_{2}}$ $(A^{-1})_{i_{2}i_{3}}$ $\cdots(A^{-1})_{i_{k}i_{1}}$. The perturbative expansion provides a straightforward method to determine the stationary point, which we will cover in detail in section 4. Figure 2 contains the Feynman diagrams for $S^{(2)},S^{(3)},$ and $S^{(4)}$. Diagrams for higher order terms simply follow this pattern.
@@
$\begin{split}-j(i)=&\sum_{k}^{M}S^{(2)}(i,k)W_{k}+\sum_{k,l}^{M}\frac{1}{2\sqrt{N}}S^{(3)}(i,k,l)(W_{k}W_{l}-W_{kl})\\ &-\sum_{k,l,m}^{M}\frac{1}{6N}S^{(4)}(i,k,l,m)\bigg[W_{kl}W_{m}+W_{km}W_{l}+W_{lm}W_{k}-W_{klm}-W_{k}W_{l}W_{m}\bigg]\end{split}$ (4)
@@
$\begin{split}0=&\sum_{k}^{M}S^{(2)}(i,k)W_{kt}+\sum_{k,l}^{M}\frac{1}{2\sqrt{N}}S^{(3)}(i,k,l)\bigg[W_{kt}W_{l}+W_{k}W_{lt}-W_{klt}\bigg]+\delta_{it}\\ &-\sum_{k,l,m}^{M}\frac{1}{6N}S^{(4)}(i,k,l,m)\bigg[(W_{klt}W_{m}+W_{kl}W_{mt}+\text{symmetric terms})-W_{klmt}\\ &\hskip 103.85237pt-(W_{kt}W_{l}W_{m}+\text{symmetric terms})\bigg].\end{split}$ (5)
Only $M(M+1)/2$ out of the $M^{2}$ equations are unique, since derivatives commute. eqs. (4) and (5), which consist of $M+M(M+1)/2$ equations in total, serve as the metrics for how successful this approximation is. Note that the sources $j(i)$ are set to 0 in eq. (4) by definition.
@@ 2.1 Approximating Higher Order Correlation Functions
By definition, we have that $W_{i}$ is the 1-pt correlation function $\langle x(i)\rangle$. There are only $M$ such terms, so these terms are manageable. Similarly, finding the 2-pt correlation function $\langle x(i)x(j)\rangle$ will help us find the connected correlation function $W_{ij}$:
$W_{ij}=W_{i}W_{j}-\langle x(i)x(j)\rangle.$ (6)
However, working with 3 and 4-pt correlation functions is a lot more computationally intensive, especially given the large number of degrees of freedom.
@@
$\begin{split}W_{3}(k,l,m)\equiv W_{klm}&=\frac{\partial^{3}W}{\partial j(k)\partial j(l)\partial j(m)}=-\frac{\partial}{\partial j(m)}(\Gamma_{kl}^{-1})=-\sum_{a}^{M}\frac{\partial s_{c}(a)}{\partial j(m)}\frac{\partial}{\partial s_{c}(a)}(\Gamma_{kl}^{-1})\\ &=\sum_{a,b,c}^{M}W_{am}W_{bk}\Gamma_{abc}W_{cl}\approx\sum_{a,b,c}^{M}W_{am}W_{bk}S^{(3)}(a,b,c)W_{cl}.\end{split}$ (8)
Above, $\Gamma^{-1}_{kl}=(\partial^{2}\Gamma/\partial s_{c}(k)\partial s_{c}(l))^{-1}=-\partial^{2}W/\partial j(k)\partial j(l)$ - a fact that directly follows from eq. (7). Similarly, $W_{klmn}$ can be expressed as:
$\begin{split}W_{4}(k,l,m,n)\equiv W_{klmn}\approx&\sum_{a,b,c}^{M}\bigg[W_{amn}W_{bk}S^{(3)}(a,b,c)W_{cl}+W_{bkn}W_{am}S^{(3)}(a,b,c)W_{cl}\\ &\hskip 28.45274pt+W_{cnl}W_{bk}S^{(3)}(a,b,c)W_{am}\bigg]\\ +&\sum_{a,b,c,d}^{M}W_{dn}W_{am}W_{bk}S^{(4)}(a,b,c,d)W_{cl}.\end{split}$ (9)
A tree diagram representation of eqs. (8) and (9) can be found in figures 3 and 4. In the derivations up until now, one might have argued that one could ignore the $1/N$ term, since we are considering the large $N$ limit. However, we included $W_{4}$ to account for its potentially large contribution given its complex structure.
@@
With these substitutions, eqs. (4) and (5) can be expressed entirely in terms of $W_{i}$ and $W_{ij}$ , drastically reducing the number of unknowns in our equations to $M+M(M+1)/2$. Since this is also the exact number of equations, either the system has one solution or none. We attempted to utilize Solve and NSolve on Mathematica to explicitly find a solution (or lack thereof). However, such a code was unable to finish running in a reasonable time, even when considering $M=4$ and ignoring the $1/N$ terms. Thus, we turned to a numerical investigation of the 1 and 2-pt correlation functions. If the numerically evaluated values satisfy the closed equations, then the approximation is successful.
@@ 3 Specifics of the Numerical Analysis
The correlation functions cannot be explicitly calculated, since $Z_{0}$ and its respective numerator diverge. Even if it were possible, such a calculation would be unreasonable for higher dimensional integrals (i.e. $M\gg 1$). This is especially troublesome for us, given that $M$ scales as a quadratic as the Euclidean lattice is taken to be a square.
@@ 3.1 Introduction to Markov Chain Monte Carlo
Markov Chain Mote Carlo (MCMC) is a numerical method that explicitly combats both of these issues budd2023monte; morningstar2007monte. The fundamental strategy of this technique is to construct a Markov Chain, the mean of which is the desired value, for each observable. A simple uniform probability distribution will also yield accurate results, but the variance of the elements would fall too slowly, especially when working with many observables morningstar2007monte. The MCMC method reduces variance by introducing two different ideas: (i) rather than uniformly choosing integration variable values, it chooses them based on a probability distribution - this is a technique called importance sampling - and (ii) element $N$ of the Markov Chain is chosen based on only element N1N-1 and through importance sampling. Incorporating both of these properties reduces how fast the variance shrinks. The mathematics that validate this numerical technique, along with error analysis, can be found in morningstar2007monte.
[…]
Figure 4: A Tree diagram illustrating the relation between $W_{4}$ and $\Gamma_{4}$, where two symmetric terms of the last diagram are omitted. Note that the $W_{3}$ term that appears in eq. (9) has already been solved for in terms of $\Gamma_{3}$. For large enough NN, $\Gamma_{4}=S^{(4)}$.
@@
Above, we used the fact that $\int[dx]e^{-S}/Z_{0}=1$, which makes $e^{-S}/Z_{0}$ a valid probability distribution $\pi(x)$ for our purposes. [^1] We choose to use the same variable names as morningstar2007monte throughout this manuscript to avoid confusion for readers who would like to expand their understanding of the numerical technique. Note that we can avoid evaluating $Z_{0}$ altogether to show this. Similarly, $W_{ij}=\langle x(i)\rangle\langle x(j)\rangle-\langle x(i)x(j)\rangle$.
A concrete methodology to effectively utilize MCMC is given by the Metropolis-Hastings algorithm. While this is certainly not the only algorithm, it allows us to construct an irreducible, aperiodic Markov Chain with positive recurrent states - conditions necessary for the success of MCMC morningstar2007monte. This algorithm revolves around updating the values of the action’s variable(s) (of a site) with the given probability distribution. If the action is multi variable - such as ours - then it is acceptable to either randomly choose a site to update or sequentially update each site. As stated prior, the updating process only takes in the current value $x(i)$ of a site to determine if the site can harbor the new value $\tilde{x}(i)$. That is, $P(\tilde{x}\leftarrow x)$ is only a function of $x,\tilde{x}$, and the probability distribution.
The general outline of the algorithm is as follows budd2023monte; morningstar2007monte:
@@ 3.2 Applying Metropolis-Hasting
Before we can work with the algorithm, we must numerically compute two quantities in our action: the discrete Laplacian and the stationary point. The perturbative expansion of the action with respect to $\sqrt{N}$ tells us that $\Phi_{0}+(A^{-1})_{ii}=0$ for fixed $i$. This generally needs to be solved numerically, as $A_{ij}$ is a function of $\Phi_{0}$. The 2D discrete Laplacian with periodic boundary conditions can be explicitly calculated using a five-point stencil on a grid like in figure 1, but this may be tedious for larger grids. A much easier alternative is using the Kronecker sum of two 1 dimensional discrete Laplacians $L_{1}$ with periodic boundary conditions. Thus, the discrete Laplacian $K_{ij}$ can be written as:
@@
Above, $\otimes$ represents a Kronecker product. All matrices on the right hand side above are $\sqrt{M}\times\sqrt{M}$ dimensional. This forces the 2D discrete Laplacian to be an $M\times M$ matrix. When constructed like this, $K_{ij}$ may not look translation invariant. That is, $K_{ij}\neq K(i-j)$. However, a simple rearranging of rows and columns can achieve this. Doing so does not change its spectrum, so the dynamics governed by the system can be studied using $K_{ij}$ from eq. (12). There is also ambiguity in the sign of this matrix, based on the stencil chosen for $L_{1}$, but changing the sign simply adds a negligible constant in the action. The stationary point also changes sign, which does not affect any calculations.
An issue with using this calculated $K_{ij}$ is the inclusion of periodic boundary conditions. This forces each row to sum to 0, making the matrix singular. When solving $\Phi_{0}+(A^{-1})_{ii}=0$, we do not find valid solutions. Some of them are complex, while the rest are not local minimums of the action (when fluctuations are turned off). [^2] This can be checked by plotting the action (with fluctuations turned off) as a function of $\Phi_{0}$. The lack of a stationary/equilibrium phase suggests that there is no stationary point here. To combat this, we redefined $K_{ij}$ by adding an identity matrix to it. Any matrix of the form $c\mathbb{1}$ would suffice, so we chose $c=1$ for simplicity. This rid the system of the singularity, allowing for a valid stationary point to exist. This extra term that we included can be thought of as a potential term.
@@
The action that we will use in the program is the closed-form variation so that we can avoid the infinite sum. However, this introduces a subtle source of error that has big ramifications. The two versions of the action are analytically the same except for two terms: the constant term and the linear term (with respect to $x(i)$). It is evident from eq. (13) that the constant terms do not matter. One may be tempted to say the same thing about the linear term, but this is not true numerically. Choosing a stationary point that makes the linear term vanish is numerically equivalent to having an extremely small (and non-zero) term. Since the linear term is magnified by the corresponding factor of $\sqrt{N}$ attached to it, the desired integrals actually grow as $\sqrt{N}$. This can be avoided if we explicitly subtract out the linear term from the action. To potentially avoid unforeseen errors, we also subtracted out the constant terms.
While the action determines the likelihood of accepting a new value at a site, it does not play a role in proposing a new value at a site. There are multiple ways to do so, but in this version of the Metropolis-Hastings algorithm, we choose a point near the current value. That is, $\tilde{x}(j)=x(j)+\delta$, where $\delta$ is uniformly and randomly chosen from an interval $\Delta,\Delta]$. There is no quantitative formula for $\Delta$, but the consensus is to choose $\Delta$ such that the program accepts new values at sites 50% – 60% of the time. This choice is to maximize efficiency and precision morningstar2007monte.
@@
Figure 5: Plot of $\langle x_{3}\rangle$ for $M=4$ and $N=100$. All $x(i)$ of each run started with the same values, which were 0 (blue/solid), 2-2 (red/dashed), and 44 (black/dot-dashed) respectively. Note that after around $\sim 300$ updates, thermalization has been achieved.
The next phase is straightforward: we continuously conduct sweeps. Here, there are two simple, computational differences: (i) all previous sets of values must be stored and (ii) after each update, the observables calculated from the site values must be stored. However, only a subset of the observables’ arrays will be elements of the Markov Chain. Before we discuss how to determine this subset, we must understand why all the measurements cannot be considered. By definition of a stochastic process, all measurements are dependent on the directly prior set of measurements. Since the updates are typically small, the statistical dependence that one set of measurements shares with the next few sets of measurements will bar the MCMC error from decreasing budd2023monte. For example, consider the following integral:
@@
The autocorrelation of the measurements can be utilized to discard these unnecessary measurements. If we were to apply Metropolis-Hasting to integrate $\int dx~f(x)p(x)$ and obtain measurements (not Markov Chain) $Y=\{f(y_{k})\}_{k=1}^{n}$, then the autocorrelation between measurements $i$ and $i+h$ is defined as:
@@
When $\rho(h)$ is small enough, we can simply skip $h$ values in between measurements to construct our Markov Chain. There is not a consensus on this threshold, but we will use $\rho(h)<0.1$ budd2023monte; morningstar2007monte. When the statistical dependence of the elements is this small, the error can simply be taken to be the Var($Y_{\text{MC}}$)/$n$, where $n$ is the length of the Markov Chain $Y_{\text{MC}}$. Thus, as $n\rightarrow\infty$, the exact value of the integral is reached.
Figure 6: Plot of $\rho(h)$ with respect to $h$ for the measurements of $\langle x(1)x(4)\rangle$ for $M=9$ and $N=40$.
To obtain these values in code, we update each site 10000 to obtain a large-enough array of measurements. While 10000 is an arbitrary number, a large-enough set of measurements is needed to ensure normal behavior from the variance. Then, we just keep calculating $\rho(h)$ until the first $h_{*}$ such that $\rho(h)<0.1$. An example of what these typically behave like can be found in figure 6. Each site $x(i)$ may have a slightly different $h_{*}$. [^3] As far as we can tell, the length of the Markov Chain does not affect $h_{*}$ drastically. We then continue the measurement phase and construct our Markov Chains with every $h_{*}$ measurements for each site. Figure 7 illustrates $\langle x^{2}(1)\rangle$, from different numbers of measurements, along with their respective errors as an example.
@@
A particularly interesting problem, to which we hope to return in the future, is the application of these methods to the Homogeneous Electron Fluid. That model has a well known first order quantum phase transition between Wigner crystal and fluid phases. Arguments have been presented that there are intermediate colloidal phases and that the transition between these and the fluid is second order. The straightforward $1/N$ approximation cannot detect the Wigner crystal and thus misses these intermediate phases as well. Our hope is that the resummation method described here, extended to larger lattices, can detect the full phase structure of the model.